For each of the following functions, respond to the given prompts. Show all work that leads to your responses.Β 

Given: 𝑓(π‘₯) = 3π‘₯^3 βˆ’ 18π‘₯^2 βˆ’ 45π‘₯ + 10Β 

a. On what interval(s), is 𝑓(π‘₯) increasing? Show all work and explain


b. At what value(s) of π‘₯ does 𝑓(π‘₯) have a relative minimum? Show all work and explain


c) On what interval(s), is 𝑓(π‘₯) decreasing and concave up? Show all work and explain


1
Expert's answer
2021-02-25T05:03:17-0500

Given: "\ud835\udc53(\ud835\udc65) = 3\ud835\udc65^3 \u2212 18\ud835\udc65^2 \u2212 45\ud835\udc65 + 10"

a. On what interval(s), is 𝑓(π‘₯) increasing?

"f'(x) > 0"

"9x^2-36x-45 = 9(x-4x-5) = 9(x-5)(x+1) > 0"

function is increasing where "(-\infin; -1)\cup (5; \infin)"

b. At what value(s) of π‘₯ does 𝑓(π‘₯) have a relative minimum?

We say that f(x) has a relative minimum at x = c if "f(x) \geqslant f(c)" for every x in some open interval around "x = c" x = -1 is relative minimum

c) On what interval(s), is 𝑓(π‘₯) decreasing and concave up?

"f'(x) < 0"

"9x^2-36x-45 = 9(x-4x-5) = 9(x-5)(x+1) < 0"

function is decreasing where "(-1;5)"

  • When the second derivative isΒ positive, the function isΒ concave upward.

"f''(x) = 18x-36 > 0"

"x > 2"

function is concave up where "(2; \infin)"